OpenStudy (anonymous):

can anyone help me with this equation the question says to solve the equation algebraically and graphically. x+1-2√(x+4)=0

OpenStudy (amoodarya):

x+1-2√(x+4)=0 then x+1=2√(x+4) (x+1)^2=(2√(x+4))^2 then simplify

OpenStudy (amoodarya):

image that i send include other solution

OpenStudy (amoodarya):

understand ?

OpenStudy (anonymous):

those it become x^2+x=1= to?

OpenStudy (anonymous):

+1 sorry

OpenStudy (amoodarya):

x^2+2x+1=4(x+4) simplify one of roots is - so its not acceptable